OpenStudy (anonymous):

Solve cube root of the quantity two y minus three equals cube root of the quantity two minus y for y. -5/3 5/3 3/5 -3/5

ganeshie8 (ganeshie8):

\( \large \sqrt[3]{2y-3} = \sqrt[3]{2-y} \)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

yes! @ganeshie8

ganeshie8 (ganeshie8):

easy then :) cube both sides

ganeshie8 (ganeshie8):

\(\large 2y-3 = 2-y\)

ganeshie8 (ganeshie8):

you can solve y now ?

OpenStudy (anonymous):

Im not sure if it is positive or negative, @ganeshie8

ganeshie8 (ganeshie8):

add y both sides

ganeshie8 (ganeshie8):

\(\large 2y-3 = 2-y\) +y +y

ganeshie8 (ganeshie8):

\(\large 3y-3 = 2\)

ganeshie8 (ganeshie8):

add 3 both sides

ganeshie8 (ganeshie8):

\(\large 3y-3 = 2\) +3 +3

ganeshie8 (ganeshie8):

\(\large 3y = 5\)

ganeshie8 (ganeshie8):

divide 3 both sides

ganeshie8 (ganeshie8):

\(\large y = 5/3\)

OpenStudy (anonymous):

Oh okay I understand now! Thanks so much! @ganeshie8

ganeshie8 (ganeshie8):

glad to hear :) you wlcme :D